Express the given equation in slope-intercept form and then state the slope “m” and the y-intercept. 3x-2y=24
Answer by ewatrrr(24785)

 
Hi,


3x-2y=24
solving for y
3x-24 = 2y
(3/2)x-12 = y
y = (3/2)x-12 , slope m = (3/2) and b = -12